OpenStudy (anonymous):

What is the domain of the function composed of f(x) = sinx g(x) = \sqrt{1 - x^2} f(g(x)) dom = ?

OpenStudy (anonymous):

make sure that \(1-x^2\geq 0\)

OpenStudy (anonymous):

do you know how to finish that?

OpenStudy (anonymous):

Wait For this one, it will be [-1,1]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wait, so then g(f(x)) = R right, because for any value of of the function, it will work? So the trick with the domains of composite functions is that the domain is always determined by the innermost function?

OpenStudy (anonymous):

well close. it is everthing in the inner function whose output is in the domain of the outer function

OpenStudy (anonymous):

but in this case since for all x \(-1\leq \sin(x)\leq 1\) the domain of \(g\circ f\) is all real numbers

OpenStudy (anonymous):

I don't really understand what you said about the composite functions. Could you give me an example to illustrate it?

OpenStudy (anonymous):

well here is an easy one. suppose you have \[f(x)=\sqrt{x}\] and \[g(x)=1-x\] now the domain of g is clearly all real numbers, and the domain of f is all numbers greater than or equal to zero but the domain of \[f\circ g\] is not all real numbers even though the domain of g is it is the set of all inputs for g are in the domain of f, that is the set where \[1-x\geq 0\] i.e \[1\geq x\]

OpenStudy (anonymous):

that was a simple example but i hope you get the idea. in this case of course \[f\circ g(x)=\sqrt{1-x}\] so the domain is \(x\leq 1\)

OpenStudy (anonymous):

So it basically has to be in the range of the outer function?

OpenStudy (anonymous):

That is, the x-values that you sub in for the innermost function?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

OH, so you can just do: range(outer most fcn) interset (domain(inner most function)

OpenStudy (anonymous):

intersect, with that upside down U

OpenStudy (anonymous):

Is that true?

OpenStudy (anonymous):

well not exactly still. it is what i wrote above the domian of the composite funciton is all numbers in the domian of the "inner function" that produce ouputs that are in the domain of outer function

OpenStudy (anonymous):

So if I wanted to the intersect of some sort, how would it look like?

OpenStudy (anonymous):

I find remembering an intersection better than the wording.

OpenStudy (anonymous):

ok but the intersection of domain of inner with what? those whose outputs are in the domain of outer. that will work

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I do not understand. Would you mind writing it out with the symbols?